# xRFM: Accurate, scalable, and interpretable feature learning models for tabular data

### v1

@@ Abstract
Inference from tabular data, collections of continuous and categorical variables organized into matrices, is a foundation for modern technology and science. Yet, in contrast to the explosive changes in the rest of AI, the best practice for these predictive tasks has been relatively unchanged and is still primarily based on variations of Gradient Boosted Decision Trees (GBDTs). Very recently, there has been renewed interest in developing state-of-the-art methods for tabular data based on recent developments in neural networks and feature learning methods. In this work, we introduce xRFM, an algorithm that combines feature learning kernel machines with a tree structure to both adapt to the local structure of the data and scale to essentially unlimited amounts of training data.
We show that compared to $31$ other methods, including recently introduced tabular foundation models (TabPFNv2) and GBDTs, xRFM achieves best performance across $100$ regression datasets and is competitive to the best methods across $200$ classification datasets outperforming GBDTs. Additionally, xRFM provides interpretability natively through the Average Gradient Outer Product. Code for xRFM (following a scikit-learn-style API) is available at: https://github.com/dmbeaglehole/xRFM.
@@ 1 Introduction
Tabular data – collections of continuous and categorical variables organized into matrices – underlies all aspects of modern commerce and science from airplane engines to biology labs to bagel shops. Yet, while Machine Learning and AI for language and vision have seen unprecedented progress, the primary methodologies of prediction from tabular data have been relatively static, dominated by variations of Gradient Boosted Decision Trees (GBDTs), such as XGBoost [7]. Nevertheless, hundreds of tabular datasets have been assembled to form extensive regression and classification benchmarks [35, 16, 12, 37, 11], and, recently, there has been renewed interest in building state-of-the-art predictive models for tabular data [18, 19, 15]. Notably, given the remarkable effectiveness of large, “foundation” models for text, there has been much excitement in developing similar models on tabular data, and recent effort has led to the development of TabPFN-v2, a foundation model for tabular data appearing in Nature [18]. Yet, despite this progress, tabular data still remains an active area for model development and building scalable, effective, and interpretable machine learning models in this domain remains an open challenge.
@@
In this work, we introduce xRFM, a tabular predictive model that combines recent advances in feature learning kernel machines with an adaptive tree structure, making it effective, scalable, and interpretable. xRFM builds upon the Recursive Feature Machine (RFM) algorithm from [27], which enabled feature learning (a form of supervised dimensionality reduction) in general machine learning models. xRFM works as follows – given training data, it first builds a binary tree structure to split data into subsets based on features relevant for prediction within each split. When splits reach a certain size (of $\leq C$ data samples), we train a leaf RFM (a hyper-parameter and compute optimized version of RFM from [28]).
@@
There are two important consequences of using a tree structure in conjunction with leaf RFMs. The first is that it enables local feature learning, meaning that xRFM can learn different features for different subsets of the data. This property is crucial for tackling hierarchical structure in tabular data (for example, one subset of features may become relevant when a specific feature takes on high value while another subset of features is relevant when the same feature takes low values). The second consequence is that it enables xRFM to scale log-linearly in the number of samples ($O(n\log n)$ given $n$ samples) during training and logarithmically ($O(\log n)$) during inference. As such, the inference time of xRFM is comparable to that of tree-based models.
@@
In practice, we show xRFM has the best performance across 100 tabular regression tasks and is competitive with state-of-the-art on 200 tabular classification tasks from the TALENT benchmark [37]. In particular, xRFM outperforms 31 other methods on these regression tasks including the GBDT and neural network variations mentioned in the opening paragraph. We show that xRFM achieves similar results on the largest datasets from the meta-test benchmark, where directly solving standard kernel machines becomes intractable on standard GPUs [19]. An additional benefit of xRFM is that it natively provides interpretability by exposing features learned and used for prediction. In particular, each leaf RFM learns features through a mathematical object known as the Average Gradient Outer Product (AGOP), whose diagonal indicates coordinates relevant for prediction and whose top eigenvectors indicate directions in data most relevant for prediction. Using examples of xRFM trained on synthetic and real data from TALENT, we show how AGOP matrices at each leaf RFM shed light on features relevant for prediction, including cases when leaf RFMs learn different features for different splits of data.
@@ 2.1 Kernel machines
A kernel machine is a non-parametric machine learning model [31, 3]. The idea behind kernel machines is that one can train a nonlinear predictive model by first transforming input data with a fixed, nonlinear feature map and then performing linear regression on the transformed data. Kernel machines make this procedure computationally tractable even for infinite dimensional feature maps by using kernel functions (inner products of feature mapped data). We describe kernel machines in the context of supervised learning below. Let $X\in\mathbb{R}^{n\times d}$ denote training inputs with ${x^{(i)}}^{T}$ denoting the example in the $i$th row of $X$ for $i\in[n]$ and $y\in\mathbb{R}^{n\times c}$ denote training labels. Let $K:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ denote a kernel function (a positive semi-definite, symmetric function). Given a regularization parameter $\lambda\geq 0$, a kernel machine trained on the data $(X,y)$ is a predictor, $\widehat{f}:\mathbb{R}^{d}\to\mathbb{R}^{c}$, of the form:
@@ 2.2 Recursive Feature Machines (RFMs)
The ability to learn task-relevant features from data is key to building effective predictors [9, 13, 1]. RFM, introduced in [27], is an algorithm that enables feature learning in general machine learning models through a mathematical object known as the Average Gradient Outer Product (AGOP). Given a predictive model $\widehat{f}:\mathbb{R}^{d}\to\mathbb{R}$ and data $S=\{x^{(1)},\ldots,x^{(n)}\}\subset\mathbb{R}^{d}$, the AGOP is defined as
@@
    A binary tree structure that splits the data into subsets, termed leaves, of a fixed maximum size (leaf size) that is independent of the number of training samples. Leaf RFMs are then trained using the data at each leaf. Splits involve stratifying data based on features relevant for prediction.
@@ 3.1 Leaf RFM
Here, we describe the changes we made to the original kernel RFM algorithm from [27] to produce leaf RFMs. The kernel RFM model was primarily built using kernels that were invariant to orthonormal transformations of data (those where $K(x,z)=K(Ux,Uz)$ for any orthogonal matrix $U$). Examples include the Gaussian kernel of the form $K(x,z)=\exp(-L\|x-z\|_{2}^{2})$ or the Laplace kernel of the form $K(x,z)=\exp(-L\|x-z\|_{2})$. Such invariance does not effectively leverage a special property of tabular data – the fact that each coordinate can be independently meaningful. This property has been hypothesized as an explanation for the superior performance of tree-based models, which compose functions of individual coordinates, over traditional neural networks on tabular datasets [16]. To account for this special structure, we introduce the following modifications to kernel-RFM:
@@
    We tune over a more general class of kernels $K_{p,q}(x,x^{\prime})=\exp(-\|x-x^{\prime}\|_{q}^{p}/L^{p})$ that are positive definite for $0<p\leq q\leq 2$ [30, Theorems 1, 5]. In particular, we use either $q=2$ (a generalized Laplace kernel) or $q=p$ (a product of 1D Laplace kernels).
@@
In addition to the above changes, we also implemented optimizations to speed up computations for kernels with $q=1$ on categorical variables and an adaptive approach for tuning bandwidth separately for data on each leaf. Additional details regarding these changes and hyperparameter search spaces for leaf RFMs on TALENT and Meta-test [19] are provided in Appendix A.
@@ 3.2 Tree-based data partitioning
We now explain how xRFM builds a binary tree structure to partition data into leaves on which leaf RFMs are trained. First, given a dataset $S$ with $n$ samples, we subsample $m$ points and train a leaf RFM (referred to as a split model) for one iteration on this subsample. The purpose of this split model is to learn a direction that can be used to partition the data into two subsets. To this end, we extract the top eigenvector, $v$, of the AGOP from the split model and create two subsets of the $n$ datapoints: $S_{1}=\{x\in S~;~v^{T}x>\text{Median}(v^{T}z~\text{for $z\in S)$}\}$ and $S_{2}=\{x\in S~;~v^{T}x\leq\text{Median}(v^{T}z~\text{for $z\in S)$}\}$. We repeat this procedure on $S_{1}$ and $S_{2}$ and their corresponding children until all leaves are less than a maximum leaf size $C$. We finally train a leaf RFM on each of these leaves. This procedure is illustrated in Fig. 1A and detailed in the ‘TreePartition’ procedure of Algorithm A.2.
@@
Figure 2: Features from xRFM trained on a synthetic example where splitting on the top AGOP direction enables xRFM to learn locally relevant features.
Note that it is possible to have split data based on unsupervised splitting procedures (e.g., those based on random projections or using top eigenvectors of PCA) [10]. A key advantage of our split approach is that it groups together data points based on the features most relevant for prediction, as is captured by the top eigenvector of the AGOP. We found that splitting in this manner outperformed splits based on unsupervised approaches, including those described above. Prior works have also split data using the random forest procedure. For example, the authors of Hollmann et al. [18] train TabPFN-v2 on the samples routed to a given leaf of a tree in a random forest. The procedure in xRFM differs as it stratifies samples by projection onto a direction rather than individual coordinates and produces only one balanced tree instead of a forest.
@@
A major advantage of xRFM over these other methods and kernel RFM itself is the ability for tree-based partitioning to learn features local to subsets of data. Here, inFig. 2, we provide an example clearly demonstrating this benefit. In this example, the function $f:\mathbb{R}^{d}\to\mathbb{R}$ depends on coordinates $x_{1},x_{3},x_{5}$ if $x_{0}>0$ and depends on coordinates $x_{9},x_{11},x_{13}$ if $x_{0}\leq 0$. xRFM learns $x_{0}$ as the initial split direction (through the AGOP of the split model trained on a random subset), then learns the relevant features for each case in separate leaf RFM models. This decomposition is to be contrasted with that of running kernel RFM from [27] in which the model would simply return that $\{x_{0},x_{1},x_{3},x_{5},x_{9},x_{11},x_{13}\}$ are all relevant features without decomposing into the two cases.
@@
Figure 3: Performance and runtime of xRFM on the TALENT benchmark (500 to 100,000 training samples). The y-axis of each plot is the shifted geometric mean of the error across all datasets in that category, while the x-axis is the average over all datasets of the training plus inference time per 1000 samples (meaning if a dataset has $n$ samples, we compute the training and inference time on the $n$ samples divide the total time by $n/1000$). Note, that times do not include the time to tune hyperparameters. Also, TabPFN-v2 cannot run on more than 1010k samples so performance and runtime on larger datasets is computed on a subsample (as was done in the TALENT benchmark). (A) nRMSE over 100 regression datasets, (B) classification error over 80 multi-class datasets, (C) classification error over 120 binary classification datasets.
We now apply xRFM to two tabular data benchmarks: the TALENT benchmark [37] and large datasets from the meta-test benchmark [19]. We use the former to evaluate xRFM on datasets of various sizes between $500$ and $100,000$ training samples. We use the latter to evaluate xRFM on larger datasets with at least $70,000$ samples (a setting in which direct linear solvers become intractable, taking more than 4040GB VRAM).
@@
When measuring performance on these datasets, we use Root Mean Square Error (RMSE) for regression tasks (as is used in TALENT), and we use classification error (1 - classification accuracy) for classification tasks. To measure performance on aggregate over a benchmark, we consider the following aggregation metrics on individual dataset performance measures: Shifted Geometric Mean, Arithmetic Mean, and Normalized Arithmetic Mean. All metrics are defined in Appendix A.
Results on the TALENT benchmark.
The TALENT benchmark consists of $300$ total datasets comparing $31$ different supervised learning algorithms [37]. Among these $31$ algorithms are strong, widely used predictive models including Gradient Boosted Decision Tree (GBDT) variants (like XGBoost, CatBoost, LightGBM), hyper-parameter optimized neural networks (RealMLP), and recent transformer-based foundation models (TabPFNv2). The TALENT benchmark consists of $100$ regression tasks, $120$ binary classification tasks, and $80$ multiclass classification tasks. The number of training samples per task varies between $500$ and $100,000$ and each task contains at least $5$ input variables.
Figure 4: Comparisons of xRFM with (A) kernel ridge regression and (B) the original RFM [28] on TALENT regression datasets (performance metric is normalized $R^{2}$ A). Each point denotes a dataset.
@@
Additionally, we validated that the overall training+inference cost for xRFM scales linearly in the number of samples, making it a highly scalable tool in practice (Fig. 5). Notably, xRFM was extremely fast on small datasets: it was at least two orders of magnitude faster than other methods when the dataset contains fewer than 3000 samples (we omitted TabPFNv2 in our comparison as it could not be run on all datasets due to restrictions on the allowable sample size, number of features, and output dimensionality). An additional comparison on datasets up to 500k samples is presented in Fig. B.1.
@@
To analyze xRFM performance on large datasets beyond those in TALENT, we considered the $17$ largest datasets in the meta-test benchmark from [19] (Table B.6). These datasets contain between 7070k and 500500k samples (7070k was near the limit, where linear solvers on 40GB VRAM GPUs become memory constrained).
We compared the performance of xRFM to other models with results reported in the literature on these $17$ datasets (Fig. 6). These other models included GBDTs (XGBoost, LightGBM, CatBoost), and neural networks (ResNet, MLP, MLP-PLR [14], RealMLP). We reported percentage improvement of each method over MLPs (following the procedure in Gorishniy et al. [15], Qu et al. [26]). Similar to the results on TALENT, we found that xRFM was best on regression tasks, and second best on classification tasks. All results are presented in Tables B.6B.7. For meta-test, we searched over slightly larger regularization parameters and slightly smaller bandwidth parameters to adapt to the much larger sample sizes of these datasets.
@@
Figure 6: Performance comparison across $17$ large datasets from meta-test (70,000-500,000 samples). (A, B) Percentage improvement over MLP error on (A) regression and (B) classification datasets. Average percentage improvement is denoted as a diamond. Points denote individual dataset results (points are jittered for visibility).

### v2

@@ Abstract
Inference from tabular data, collections of continuous and categorical variables organized into matrices, is a foundation for modern technology and science. Yet, in contrast to the dramatic developments in the rest of AI, the best practice for these predictive tasks has been relatively unchanged and is still primarily based on variations of Gradient Boosted Decision Trees (GBDTs). Very recently, there has been renewed interest in developing state-of-the-art methods for tabular data based on recent developments in neural networks and feature learning methods. In this work, we introduce xRFM, an algorithm that combines feature learning kernel machines with a tree structure to both adapt to the local structure of the data and scale to essentially unlimited amounts of training data.
On the TALENT benchmark, we show that compared to $31$ other methods, including recently introduced tabular foundation models (TabPFNv2) and GBDTs, xRFM achieves best performance across $100$ regression datasets and is competitive to the best methods across $200$ classification datasets outperforming GBDTs. Additionally, xRFM provides interpretability natively through the Average Gradient Outer Product. Code for xRFM (following a scikit-learn-style API) is available at: https://github.com/dmbeaglehole/xRFM.
@@ 1 Introduction
Tabular data – collections of continuous and categorical variables organized into matrices – underlies all aspects of modern commerce and science from airplane engines to biology labs to bagel shops. Yet, while methods for Machine Learning and AI in language and vision have seen unprecedented progress, the primary methodologies of prediction from tabular data have been relatively static, dominated by variations of Gradient Boosted Decision Trees (GBDTs), such as XGBoost [8]. Nevertheless, hundreds of tabular datasets have been assembled to form extensive regression and classification benchmarks [6, 18, 14, 38, 13], and, recently, there has been renewed interest in building state-of-the-art predictive models for tabular data [20, 21, 17]. Notably, given the remarkable effectiveness of large, “foundation” models for text, there has been much excitement in developing similar models on tabular data, and recent effort has led to the development of TabPFN-v2, a foundation model for tabular data appearing in Nature [20]. Yet, despite this progress, building scalable, effective, and interpretable machine learning models in this domain remains an open challenge.
@@
In this work, we introduce xRFM, a tabular predictive model that combines recent advances in feature learning kernel machines with an adaptive tree structure, making it highly accurate, scalable, and interpretable. xRFM builds upon the Recursive Feature Machine (RFM) algorithm from [29], which enabled feature learning (a form of supervised dimensionality reduction) in general machine learning models. xRFM works as follows: Given training data, it first builds a binary tree structure to split data into subsets based on features relevant for prediction within each split. When splits reach a certain size (of $\leq C$ data samples), we train a leaf RFM (a hyper-parameter and compute optimized version of the original RFM).
@@
There are two important consequences of using a tree structure in conjunction with leaf RFMs. The first is that it enables local feature learning, meaning that xRFM can learn different features for different subsets of the data. This property is crucial for tackling hierarchical structure in tabular data (for example, one set of features may be relevant when a specific feature takes on high value while another set of features is relevant when the same feature takes low values). The second consequence is that it enables xRFM to scale nearly linearly in the number of samples ($O(n\log n)$ given $n$ samples) during training and logarithmically ($O(\log n)$) during inference. As such, the inference time of xRFM is comparable to that of tree-based models.
@@
In practice, we show xRFM has the best performance across 100 tabular regression tasks and is competitive with state-of-the-art on 200 tabular classification tasks from the TALENT benchmark [38]. In particular, xRFM outperforms 31 other methods on these regression tasks including the GBDT and neural network variations mentioned in the opening paragraph. On the TabArena-Lite benchmark [13], we show xRFM achieves one of the best tradeoffs (is on the empirical Pareto frontier) between performance and inference time among all methods in regression and is again competitive on classification. We show that xRFM achieves similar results on the largest datasets from the meta-test benchmark [21], where directly solving standard kernel machines becomes intractable on standard GPUs. An additional benefit of xRFM is that it natively provides interpretability by exposing features learned and used for prediction. In particular, each leaf RFM learns features through a mathematical object known as the Average Gradient Outer Product (AGOP), whose diagonal indicates coordinates relevant for prediction and whose top eigenvectors indicate directions in data most relevant for prediction. Using examples of xRFM trained on synthetic and real data, we show how AGOP matrices at each leaf RFM shed light on features relevant for prediction, including cases when leaf RFMs learn different features for different splits of data.
@@ 2.1 Kernel machines
A kernel machine is a non-parametric machine learning model [33, 3]. The idea behind kernel machines is that a nonlinear predictive model can be trained by first transforming input data with a fixed, nonlinear feature map and then performing linear regression on the transformed data. Kernel machines make this procedure computationally tractable even for infinite dimensional feature maps by using explicitly defined kernel functions (inner products of feature mapped data). We describe kernel machines in the context of supervised learning below. Let $X\in\mathbb{R}^{n\times d}$ denote training inputs with ${x^{(i)}}^{T}$ denoting the example in the $i$th row of $X$ for $i\in[n]$ and $y\in\mathbb{R}^{n\times c}$ denote training labels. Let $K:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ denote a kernel function (a positive semi-definite, symmetric function). Given a regularization parameter $\lambda\geq 0$, a kernel machine trained on the data $(X,y)$ is a predictor, $\widehat{f}:\mathbb{R}^{d}\to\mathbb{R}^{c}$, of the form:
@@ 2.2 Recursive Feature Machines (RFMs)
The ability to learn task-relevant features from data is key to building effective predictors [10, 15, 1]. RFM, introduced in Radhakrishnan et al. [29], is an algorithm that enables feature learning in general machine learning models through a mathematical object known as the Average Gradient Outer Product (AGOP). Given a predictive model $\widehat{f}:\mathbb{R}^{d}\to\mathbb{R}$ and data $S=\{x^{(1)},\ldots,x^{(n)}\}\subset\mathbb{R}^{d}$, the AGOP is defined as
@@
    A binary tree that splits the data into subsets, termed leaves, of a maximum size (leaf size) that is independent of the number of training samples. Leaf RFMs are then trained on the subset at each leaf. The splits stratify data based on features relevant for prediction.
@@ 3.1 Leaf RFM
Here, we describe the changes we made to the original kernel RFM algorithm from Radhakrishnan et al. [29] to produce leaf RFMs. The kernel RFM model was primarily built using kernels that were invariant to orthonormal transformations of data (those where $K(x,z)=K(Ux,Uz)$ for any orthogonal matrix $U$). Examples include the Gaussian kernel of the form $K(x,z)=\exp(-L\|x-z\|_{2}^{2})$ or the Laplace kernel of the form $K(x,z)=\exp(-L\|x-z\|_{2})$. Such invariance does not effectively leverage a special property of tabular data – the fact that each coordinate can be independently meaningful. This property has been hypothesized as an explanation for the superior performance of tree-based models, which compose functions of individual coordinates, over traditional neural networks on tabular datasets [18]. To account for this special structure, we introduce the following modifications to kernel-RFM:
Figure 2: Training xRFM on synthetic data where splitting on the top AGOP direction enables xRFM to learn locally relevant features.
@@
    We tune over a more general class of kernels $K_{p,q}(x,x^{\prime})=\exp(-\|x-x^{\prime}\|_{p}^{q}/L^{q})$ that are positive definite for $0<q\leq p\leq 2$ [32, Theorems 1, 5].
@@
In addition to the above changes, we also implemented optimizations to speed up computations for categorical variables and an adaptive approach for tuning bandwidth separately for data on each leaf. Additional details regarding these changes and hyperparameter search spaces for leaf RFMs on TALENT, TabArena-Lite, and meta-test are provided in Appendix A.
@@ 3.2 Tree-based data partitioning
We now explain how xRFM builds a binary tree structure to partition data into leaves on which leaf RFMs are trained. First, given a dataset $S$ with $n$ samples, we subsample $m$ points and train a leaf RFM (referred to as a split model) for one iteration on this subsample. The split model serves to learn a direction that can be used to partition the data into two subsets. To this end, we extract the top eigenvector, $v$, of the AGOP from the split model and create two subsets of the $n$ datapoints: $S_{1}=\{x\in S~;~v^{T}x>\text{Median}(v^{T}z~\text{for $z\in S)$}\}$ and $S_{2}=\{x\in S~;~v^{T}x\leq\text{Median}(v^{T}z~\text{for $z\in S)$}\}$. We repeat this procedure on $S_{1}$ and $S_{2}$ and their corresponding children until all leaves are less than a maximum leaf size $C$. We finally train a leaf RFM on each of these leaves. This procedure is illustrated in Fig. 1A and detailed in the ‘TreePartition’ procedure of Algorithm A.2.
@@
Note that it is possible to have split data based on unsupervised splitting procedures (e.g., those based on random projections or using top eigenvectors of PCA) [11]. A key advantage of our split approach is that it groups together data points based on the features most relevant for prediction, as is captured by the top eigenvector of the AGOP. We found that splitting in this manner outperformed splits based on unsupervised approaches, including those described above. Prior works have also split data using the random forest procedure. For example, Hollmann et al. [20] train TabPFN-v2 on the samples routed to a given leaf of a tree in a random forest. The procedure in xRFM differs as it stratifies samples by projection onto a direction rather than individual coordinates and produces only one balanced tree instead of a forest.
Figure 3: Performance and runtime of xRFM on the TALENT (Plots A-C) and TabArena-Lite benchmarks (Plots D-F). The y-axes of plots A-C are the shifted geometric mean of the error across all datasets in that category, while the x-axes are the average over all datasets of the training plus inference time per 1000 samples for just the best hyperparameter configuration (meaning if a dataset has $n$ samples, we compute the training and inference time on the $n$ samples divide the total time by $n/1000$). The y-axes in plots D-F are Elo, the main metric used in TabArena, and reflect the relative win-rate of each method, while the x-axes are the median inference time per 1000 total samples. The TabArena-Lite plots additionally show the default methods and the compute Pareto tradeoff curve for inference time versus Elo. The TALENT plots are (A) nRMSE over 100 regression datasets, (B) classification error over 80 multi-class datasets, (C) classification error over 120 binary classification datasets. The TabArena-Lite plots are (D) regression, (E) multi-class, and (F) binary datasets.
@@
A major advantage of xRFM over these other methods and kernel RFM itself is the ability for tree-based partitioning to learn features local to subsets of data. Here, in Fig. 2, we provide an example clearly demonstrating this benefit. In this example, the function $f:\mathbb{R}^{d}\to\mathbb{R}$ depends on coordinates $x_{1},x_{3},x_{5}$ if $x_{0}>0$ and depends on coordinates $x_{9},x_{11},x_{13}$ if $x_{0}\leq 0$. xRFM learns $x_{0}$ as the initial split direction (through the AGOP of the split model trained on a random subset), then learns the relevant features for each case in separate leaf RFM models. In contrast, if running kernel RFM from Radhakrishnan et al. [29], the model would simply return that $\{x_{0},x_{1},x_{3},x_{5},x_{9},x_{11},x_{13}\}$ are all relevant features without decomposing into the two cases.
@@
[…]
Figure 4: Comparisons of xRFM with (A) kernel ridge regression and (B) the original RFM [30] on TALENT regression datasets (metric is normalized $R^{2}$, see Appendix A). Each point is a dataset.
@@
When measuring performance on these datasets, we use Root Mean Square Error (RMSE) for regression tasks (as is used in TALENT), and we use classification error ($1$ - classification accuracy) for classification tasks. For TabArena-Lite, binary classification performance for individual datasets are measured by AUROC, while multi-class datasets are evaluted with log-loss. To measure performance on aggregate over TALENT, we consider the following aggregation metrics on individual dataset performance measures: Shifted Geometric Mean, Arithmetic Mean, and Normalized Arithmetic Mean. TabArena-Lite uses Elo [12], a rating system that calculates the relative skill levels of each method based on the probability of winning when pairs of methods are compared on individual datasets (see Erickson et al. [13] for how Elo is estimated in this benchmark). All other metrics are defined in Appendix A.
Results on the TALENT and TabArena-Lite benchmarks.
The TALENT benchmark consists of $300$ total datasets comparing $31$ different supervised learning algorithms [38]. Among these $31$ algorithms are strong, widely used predictive models including Gradient Boosted Decision Tree (GBDT) variants (like XGBoost, CatBoost, LightGBM), hyper-parameter optimized neural networks (RealMLP), and recent transformer-based foundation models (TabPFN-v2). The TALENT benchmark consists of $100$ regression tasks, $120$ binary classification tasks, and $80$ multiclass classification tasks. The number of training samples per task varies between $500$ and $100,000$ and each task contains at least $5$ input variables. TabArena-Lite contains $51$ total datasets ($13$ regression and $38$ classification) comparing 15 different supervised learning algorithms including many of those in TALENT. The datasets in TabArena-Lite vary in sample size from $700$ to $150,000$.
@@
Fig. 5 shows that the overall training+inference cost of xRFM initially scales super-linearly. However, xRFM is still reasonably fast at 60K samples, after which the log-linear scaling thanks to the tree-based data partitioning kicks in (Fig. B.1). Notably, xRFM is extremely fast on small datasets: it is at least two orders of magnitude faster than other methods when the dataset contains fewer than 3000 samples (we omit TabPFN-v2 in our comparison as it could not be run on all datasets due to restrictions on the allowable sample size, number of features, and output dimensionality).
@@
To analyze xRFM performance on large datasets beyond those in TALENT and TabArena, we consider the $17$ largest datasets in the meta-test benchmark from [21] (Table B.6).
We compare the performance of xRFM to other models with results reported in the literature on these $17$ datasets (Fig. 6). These other models include GBDTs (XGBoost, LightGBM, CatBoost), and neural networks (ResNet, MLP, MLP-PLR [16], RealMLP). We report the percentage improvement of each method over MLPs, following the procedure in Gorishniy et al. [17], Qu et al. [28]. Similar to the results on TALENT, xRFM is best on regression tasks, and second best on classification tasks. All results are presented in Tables B.6B.7. For meta-test, we search over slightly larger regularization parameters and slightly smaller bandwidth parameters to adapt to the much larger sample sizes of these datasets.
@@
Figure 6: Performance comparison across the $17$ large datasets from meta-test (70,000-500,000 samples). (A, B) Percentage improvement over MLP error on (A) regression and (B) classification datasets. Average percentage improvement is denoted as a diamond. Points denote individual dataset results (points are jittered for visibility).